OpenStudy (anonymous):

How many molecules of H2O are there in 1.0 g of H2O?

OpenStudy (anonymous):

the molar mass of H2O=18 no. of moles = mass/molar mass =1/18 mole no. of molecules = no. of moles*avogadro's number avogaro's number= 6.022*10^23 so no. of molecules =1/18 *6.022*10^23